OpenStudy (anonymous):

Electrical energy (E) is measured in kilowatt hours (kwh) using the formula E = P · t, where P is the power in kilowatts and t is the time in hours. If 84 kwh of electrical energy are created over a period of 14 hours, how many kilowatts of power are used each hour? 6 70 98 1,176

OpenStudy (zpupster):

your formula E=P*t we know E = 84kwh and t =14 h divide E/t= P kilowatts of power